# MMICT: Boosting Multi-Modal Fine-Tuning with In-Context Examples

### v2

@@ Abstract
Although In-Context Learning (ICL) brings remarkable performance gains to Large Language Models (LLMs), the improvements remain lower than fine-tuning on downstream tasks. This paper introduces Multi-Modal In-Context Tuning (MMICT), a novel multi-modal fine-tuning paradigm that boosts multi-modal fine-tuning by fully leveraging the promising ICL capability of multi-modal LLMs (MM-LLMs). We propose the Multi-Modal Hub (M-Hub), a unified module that captures various multi-modal features according to different inputs and objectives. Based on M-Hub, MMICT enables MM-LLMs to learn from in-context visual-guided textual features and subsequently generate outputs conditioned on the textual-guided visual features. Moreover, leveraging the flexibility of M-Hub, we design a variety of in-context demonstrations. Extensive experiments on a diverse range of downstream multi-modal tasks demonstrate that MMICT significantly outperforms traditional fine-tuning strategy and the vanilla ICT method that directly takes the concatenation of all information from different modalities as input.
1 Introduction
@@
Recently, a great number of works on large-scale language models (LLMs) (Zhao et al. 2023) have sprung up, propelling the evolution of human-like artificial intelligence. By escalating the model size, for instance, from 1 billion (GPT 1) to 175 billion parameters (GPT 3) or more, LLMs can demonstrate extraordinary proficiency in comprehending human language. Many researchers attempt to further augment the text-based LLMs by incorporating additional modalities (e.g., image, and video), leading to the creation of multi-modal LLMs (MM-LLMs). Representative works include but not limited to KOSMOS-1 (Huang et al. 2023), Flamingo (Alayrac et al. 2022) and GPT-4 (OpenAI 2023).
@@
With the prosperous development, LLMs have shown a capacity for in-context learning (ICL) (Dong et al. 2023), which involves learning and prediction solely based on a few examples in the context and does not update any parameters of LLMs. For instance, in the field of MM-LLMs, Flamingo capitalizes on the interleaved multi-modal data to enhance its multi-modal in-context learning capabilities. Building upon OpenFlamingo (Awadalla et al. 2023), the open-source version of Flamingo, Otter (Li et al. 2023b) is capable of executing new instructions with a few in-context learning examples using multi-modal in-context instruction tuning. Additionally, some studies (Zhao et al. 2021; Lu et al. 2022) verify that various demonstration factors (e.g., demonstration format and demonstration order) heavily affect the performance of ICL.
@@
Although ICL can bring remarkable performance gains to MM-LLMs, the improvements still lag behind fine-tuning on training data for downstream tasks (Alayrac et al. 2022; Chen et al. 2022b). For example, on the VQAv2 task, Flamingo achieves 63.1% accuracy with 4 demonstration examples while the accuracy after fine-tuning is 82.0%. The observation inspires us: Can we combine the two learning paradigms by leveraging ICL to further enhance the fine-tuning performance on downstream multi-modal tasks?
@@
MMICT is built based on BLIP-2 (Li et al. 2023c). BLIP-2 adopts a traditional fine-tuning strategy with only query inputs, and exclusively utilizes the cross-modal pre-trained Qformer to extract visual features. To better capture the multi-modal features within a unified model architecture, we design the Multi-Modal Hub (M-Hub) used in MMICT. Different from Qformer, M-Hub can produce either uni-modal features or multi-modal features that fuse information from different modalities. Considering that different demonstration factors (e.g., feature extraction strategy, sampling number for demonstrations and sampling strategies for demonstrations) may heavily affect the performance, we design various variants of in-context demonstrations by leveraging the flexibility of M-Hub.
@@
2 Related Work
Figure 1: Overview of MMICT. M-Hub can output both visual-guided textual features (upper left green part) and instruction-guided visual features (upper right orange part). MMICT learns from visually-guided textual features derived from demonstration examples and generates outputs based on instruction-guided visual features obtained from input queries.
2.1 Multi-Modal Large Language Models
@@
Pioneering studies like VisualGPT (Chen et al. 2022a) and Frozen (Tsimpoukelli et al. 2021) have demonstrated the benefits of employing a pre-trained language model as a vision-language model decoder. Flamingo (Alayrac et al. 2022) is proposed to align a pre-trained vision encoder and language model using the gated cross-attention mechanism. It is trained on billions of image-text pairs, showcasing impressive in-context few-shot learning capabilities. BLIP-2 (Li et al. 2023c) introduces a Q-Former to efficiently align visual features with the language model. GPT-4 (OpenAI 2023) shows more powerful visual understanding and reasoning abilities after pre-training on a vast collection of aligned image-text data. To empower LLMs with the ability of video understanding, a multi-branch cross-modal pre-training framework Video-LLaMA (Zhang, Li, and Bing 2023) is proposed to achieve both vision-language alignment and audio-language alignment by connecting the LLM to off-the-shelf uni-modal pre-trained models.
@@
In summary, there are three mainstream methods that combine visual encoder and LLM into MM-LLM: 1) The addition of layers and utilization of cross-attention in the LLM, exemplified by Flamingo and Otter, which incurs high computational costs. 2) The incorporation of learnable layers subsequent to the visual encoder, facilitating the conversion of features to enhance the comprehension of LLM, as demonstrated in the BLIP-2 and MiniGPT4 (Zhu et al. 2023), which offers a simpler and less hardware-intensive way. 3) The inclusion of supplementary learnable parameters in the LLM, as seen in the LLaMA-Adapter (Gao et al. 2023), yields a faster training speed albeit with marginally lower performance in comparison to alternative approaches. Considering the pros and cons of the aforementioned methods, we opt for BLIP-2 as the base of MMICT.
2.2 In-Context Learning
@@
In-context learning involves learning based on only a few examples in the form of demonstration. Essentially, it estimates the likelihood of the potential answer conditioned on the demonstration using a well-trained language model. For multi-modal tasks, Flamingo (Alayrac et al. 2022) capitalizes on the interleaved multi-modal data to enhance its few-shot ICL capabilities. Moreover, the paradigm of generating query text conditioned on in-context examples ensures its ICL capacity during the inference phase. Building upon OpenFlamingo, Otter (Li et al. 2023b) introduces the in-context instruction tuning paradigm for multi-modal models. However, they do not capture cross-modality information and only use uni-modal features from different modalities as demonstrations. The multi-modal features are modeled in the cross-attention module of the LLM. Differently, MMICT models cross-modality information to guide the construction of the multi-modal demonstrations.
@@
As demonstrations play a vital role in ICL, many works study demonstration designing strategies. For instance, in natural language processing, several works aim to select good examples for ICL through unsupervised methods based on pre-defined metrics (Li and Qiu 2023) or supervised methods (Wang, Zhu, and Wang 2023). For tasks requiring complex reasoning (e.g., math word problems and commonsense reasoning), some works design better demonstrations for ICL by describing tasks with the instruction (Wang et al. 2023) and adding intermediate reasoning steps (Xu et al. 2023a). Differently, MMICT proposes a better feature extraction strategy for demonstrations to avoid information redundancy and exhibits robustness against different demonstration surfaces.
3 Our Method
@@
3.1 Multi-Modal In-Context Tuning
@@
In this study, we feed paired features from different modalities into the Multi-modal Hub (M-Hub, see Sec. 3.2) to obtain multi-modal fused features. Moreover, considering the modality gap between vision and text, we retain the visual-guided textual features as demonstrations. To elucidate the detailed formulation of MMICT, let us take video captioning as an example. As depicted in Fig. 1, given a video-instruction pair $\{v_{pred},t_{ins}\}$ (upper right orange part in Fig. 1) accompanied with other pairs $\{v_{1},t_{1},...,v_{n_{e}},t_{n_{e}}\}$ (upper left green part in Fig. 1) that are randomly selected as demonstration examples, where $\{v_{*},t_{*}\}$ denotes a video clip and the corresponding text, and $n_{e}$ is the number of demonstration examples. The model needs to predict the label $y$ (e.g., the caption text “a cat is laying down washing his face”) paired with $v_{pred}$ according to the in-context information. The frozen image encoder (e.g. EVA (Fang et al. 2022)) takes as input the video clips $\{v_{1},...,v_{n_{e}},v_{pred}\}$ and outputs the corresponding encoded visual features $\{\mathbf{Z}^{v}_{1},...,\mathbf{Z}^{v}_{n_{e}},\mathbf{Z}^{v}_{pred}\}$. We first feed these features and their paired text into M-Hub, and then pass them to a share-weight fully-connected network to extract visual-guided textual demonstration features and textual-guided visual features:
$\mathbf{\hat{T}}^{c}_{k}=\text{FC}\big{(}\mathcal{G}(\mathbf{Z}_{k}^{v},t_{k})\big{)}$ (1)
$\mathbf{\hat{V}}^{d}_{pred}=\text{FC}\big{(}\mathcal{G}(\mathbf{Z}_{pred}^{v},t_{ins})\big{)}$ (2)
@@
$\mathbf{C}=\big{[}\mathbf{\hat{T}}^{c}_{1},\left\langle\text{EOC}\right\rangle,...,\mathbf{\hat{T}}^{c}_{n_{e}},\left\langle\text{EOC}\right\rangle,\mathbf{\hat{V}}^{d}_{pred}\big{]}$ (3)
@@
3.2 Multi-Modal Hub
Owing to the huge model size of LLMs and vision foundation models, training their parameters for multi-modal tasks proves to be challenging. As a result, many researchers endeavor to incorporate a comparatively lightweight and trainable Visual Prompt Generator (VPG) between them while maintaining their fully frozen state (Zhang et al. 2023). Among these efforts, the BLIP-style multi-modal pre-training approach effectively connects LLMs to vision foundation models via the Q-former. However, previous works (Zhu et al. 2023; Li et al. 2023c) solely utilize this module to extract uni-modal visual features for LLMs after pre-training, thereby overlooking the benefits of multi-modal pre-training.
@@ 3.2 Multi-Modal Hub
$\mathbf{P}_{h},\mathbf{R}_{h}=\text{Self-Attention}\big{(}[\mathbf{\hat{V}}^{d}_{h-1},\mathbf{\hat{T}}^{c}_{h-1}]\big{)}$ (4)
@@ 3.3 Training Objective
$\mathcal{L}=-\sum_{y\in Y}\sum_{t=1}^{\left|y\right|}log\big{(}p(y_{t}|\mathbf{C},t_{ins},y_{1},...,y_{t-1})\big{)}$ (5)
@@ 3.4 Inference
$\hat{y}=\text{LLM}\big{(}[\mathbf{\hat{V}}^{d}_{pred},\mathcal{T}(t_{ins})]\big{)}$ (6)
@@
Table 1: Instruction templates used for captioning tasks.
@@
4 Experiments
Table 2: Performance of all methods. We mainly compare our MMICT with two baselines. The performance of SOTA methods is denoted in gray. The best results across different types of MM-LLMs are shown in bold.
Caption VQA VideoQA
COCO MSR-VTT MSVD VQAv2 MSR-VTT MSVD
Method LLM B@4 C B@4 C B@4 C Acc Acc Acc
VLAB (He et al. 2023) - 152.5 54.6 74.9 79.3 179.8 - 49.6 61.0
VAST (Chen et al. 2023a) - 149.0 - 78.0 - - - 50.1 60.2
mPLUG-2 (Xu et al. 2023b) 41.6 137.7 57.8 80.3 75.0 165.8 - 48.0 58.1
VanillaFT FlanT5 43.4 144.8 51.3 74.7 73.0 174.3 69.6 43.4 63.9
MMICT 43.6 145.7 51.6 74.8 76.5 179.1 72.3 45.6 66.7
VanillaFT OPT 44.0 144.3 47.6 69.9 80.3 177.1 56.9 42.6 56.9
MMICT 43.9 145.5 52.0 71.4 80.4 180.4 73.0 46.0 66.3
4.1 Implementation Details
[…]
4.2 Evaluation Tasks and Datasets

### v3

@@ Abstract.
Although In-Context Learning (ICL) brings remarkable performance gains to Large Language Models (LLMs), the improvements remain lower than fine-tuning on downstream tasks. This paper introduces Multi-Modal In-Context Tuning (MMICT), a novel multi-modal fine-tuning paradigm that boosts multi-modal fine-tuning by fully leveraging the promising ICL capability of multi-modal LLMs (MM-LLMs). We propose the Multi-Modal Hub (M-Hub), a unified module that captures various multi-modal features according to different inputs and objectives. Based on M-Hub, MMICT enables MM-LLMs to learn from in-context visual-guided textual features and subsequently generate outputs conditioned on the textual-guided visual features. Moreover, leveraging the flexibility of M-Hub, we design a variety of in-context demonstrations. Extensive experiments on a diverse range of downstream multi-modal tasks demonstrate that MMICT significantly outperforms traditional fine-tuning strategy and the vanilla ICT method that directly takes the concatenation of all information from different modalities as input. Our implementation is available at: https://github.com/KDEGroup/MMICT.
Multi-Modal Alignment, Text Generation, In-Context Tuning
copyright: acmlicensedjournalyear: 2024journal: ToMMccs: Computing methodologies Natural language generationccs: Computing methodologies Computer vision tasks
1. Introduction
@@
Recently, a great number of works on large-scale language models (LLMs) (Zhao et al., 2023) have sprung up, propelling the evolution of human-like artificial intelligence. By escalating the model size, for instance, from 1 billion (GPT 1) to 175 billion parameters (GPT 3) or more, LLMs can demonstrate extraordinary proficiency in comprehending human language. Many researchers attempt to further augment the text-based LLMs by incorporating additional modalities (e.g., image, and video), leading to the creation of multi-modal LLMs (MM-LLMs). Representative works include but not limited to KOSMOS-1 (Huang et al., 2023), Flamingo (Alayrac et al., 2022) and GPT-4 (OpenAI, 2023).
@@
With the prosperous development, LLMs have shown a capacity for in-context learning (ICL) (Dong et al., 2023a), which involves learning and prediction solely based on a few examples in the context and does not update any parameters of LLMs. For instance, in the field of MM-LLMs, Flamingo capitalizes on the interleaved multi-modal data to enhance its multi-modal in-context learning capabilities. Building upon OpenFlamingo (Awadalla et al., 2023), the open-source version of Flamingo, Otter (Li et al., 2023b) is capable of executing new instructions with a few in-context learning examples using multi-modal in-context instruction tuning. Additionally, some studies (Zhao et al., 2021; Lu et al., 2022) verify that various demonstration factors (e.g., demonstration format and demonstration order) heavily affect the performance of ICL.
@@
Although ICL can bring remarkable performance gains to MM-LLMs, the improvements still lag behind fine-tuning on training data for downstream tasks (Alayrac et al., 2022; Chen et al., 2022a). For example, on the VQAv2 task, Flamingo achieves 63.1% accuracy with 4 demonstration examples while the accuracy after fine-tuning is 82.0%. The observation inspires us: Can we combine the two learning paradigms by leveraging ICL to further enhance the fine-tuning performance on downstream multi-modal tasks?
@@
MMICT is built based on BLIP-2 (Li et al., 2023a). BLIP-2 adopts a traditional fine-tuning strategy with only query inputs, and exclusively utilizes the cross-modal pre-trained Qformer to extract visual features. To better capture the multi-modal features within a unified model architecture, we design the Multi-Modal Hub (M-Hub) used in MMICT. Different from Qformer, M-Hub can produce either uni-modal features or multi-modal features that fuse information from different modalities. Considering that different demonstration factors (e.g., feature extraction strategy, sampling number for demonstrations and sampling strategies for demonstrations) may heavily affect the performance, we design various variants of in-context demonstrations by leveraging the flexibility of M-Hub.
@@
2. Related Work
2.1. Multi-Modal Large Language Models
@@
Pioneering studies like VisualGPT (Chen et al., 2022b) and Frozen (Tsimpoukelli et al., 2021) have demonstrated the benefits of employing a pre-trained language model as a vision-language model decoder. Flamingo (Alayrac et al., 2022) is proposed to align a pre-trained vision encoder and language model using the gated cross-attention mechanism. It is trained on billions of image-text pairs, showcasing impressive in-context few-shot learning capabilities. BLIP-2 (Li et al., 2023a) introduces a Q-Former to efficiently align visual features with the language model. GPT-4 (OpenAI, 2023) shows more powerful visual understanding and reasoning abilities after pre-training on a vast collection of aligned image-text data. To empower LLMs with the ability of video understanding, a multi-branch cross-modal pre-training framework Video-LLaMA (Zhang et al., 2023b) is proposed to achieve both vision-language alignment and audio-language alignment by connecting the LLM to off-the-shelf uni-modal pre-trained models. MoE-LLaVA (Lin et al., 2024) constructs a MoE-based sparse LVLM architecture with an outrageous number of parameters but a constant computational cost. Gemini family (Anil et al., 2023) exhibits remarkable capabilities across image, audio, video, and text understanding even on memory-constrained devices.
@@
In summary, there are four mainstream methods that combine visual encoder and LLM into MM-LLM: 1) The addition of extra modules in the LLM that enable deep interaction and fusion, exemplified by Otter (Li et al., 2023b) and CogVLM (Wang et al., 2023b), which incurs high computational costs. 2) The incorporation of learnable query tokens to extract information in a query-based manner, facilitating the conversion of features to enhance the comprehension of LLM, as demonstrated in the BLIP-2 (Li et al., 2023a), X-LLM (Chen et al., 2023a) and MiniGPT4 (Zhu et al., 2023), which compress visual tokens into a smaller number of representation vectors. 3) The inclusion of supplementary learnable parameters in the LLM, as seen in the LLaMA-Adapter (Gao et al., 2023), yields a faster training speed albeit with marginally lower performance in comparison to alternative approaches. 4) The simple usage of a MLP-based interface to bridge the modality gap. For example, LLaVA series adopts one/two linear MLP (Liu et al., 2023a, b) to project visual tokens and align the feature dimension with word embeddings. Despite their simplicity, the large amount of prefix visual tokens increases the context length in MM-LLM. Considering the pros and cons of the aforementioned methods, we opt for BLIP-2 as the base of MMICT.
2.2. In-Context Learning
@@
In-context learning involves learning based on only a few examples in the form of demonstration. Essentially, it estimates the likelihood of the potential answer conditioned on the demonstration using a well-trained language model. For multi-modal tasks, Flamingo (Alayrac et al., 2022) capitalizes on the interleaved multi-modal data to enhance its few-shot ICL capabilities. Moreover, the paradigm of generating query text conditioned on in-context examples ensures its ICL capacity during the inference phase. Building upon OpenFlamingo, Otter (Li et al., 2023b) introduces the in-context instruction tuning paradigm for multi-modal models. However, they do not capture cross-modality information and only use uni-modal features from different modalities as demonstrations. The multi-modal features are modeled in the cross-attention module of the LLM. Differently, MMICT models cross-modality information to guide the construction of the multi-modal demonstrations. More detailed comparisons are provided in Sec. 6.
@@
As demonstrations play a vital role in ICL, many works study demonstration designing strategies. For instance, in natural language processing, several works aim to select good examples for ICL through unsupervised methods based on pre-defined metrics (Li and Qiu, 2023) or supervised methods (Wang et al., 2023c). For tasks requiring complex reasoning (e.g., math word problems and commonsense reasoning), some works design better demonstrations for ICL by describing tasks with the instruction (Wang et al., 2023a) and adding intermediate reasoning steps (Xu et al., 2023a). Differently, MMICT proposes a better feature extraction strategy for demonstrations to avoid information redundancy and exhibits robustness against different demonstration surfaces.
2.3. Multi-modal Alignment
Multi-modal tasks necessitate a model that adeptly translates visual scenes into language descriptions. Currently popular approaches (Shi et al., 2023; Li et al., 2022; Dong et al., 2023b) leverage an encoder-decoder architecture and achieve promising results. An optimal multi-modal model embodies the trifecta of visual comprehension, visual-text alignment, and textual generation. Initially, a diverse array of feature extractors is employed to procure visual representations, fundamental for basic visual understanding. Subsequently, these visual representations are transformed into a hidden vector $h$, which shares a semantic space with the language representations’ hidden vector, thus bridging the gaps between vision and language. In essence, this alignment occurs in a shared semantic space, uniting the visual modality (e.g., video or image) with the textual modality. Ultimately, the language representations undergo decoding to produce the definitive descriptions. The pivotal step in this multi-modal model is the visual-text alignment, mapping visual representations to the language domain. This alignment is crucial because the language network (such as LSTM or Transformer) within the decoder can unleash its full text generation prowess when the input representation resonates with the language domain. However, existing techniques tend to concentrate either on video comprehension (Peng et al., 2024; Tang et al., 2019) or image understanding (Wang et al., 2018; Jiang et al., 2021). Thanks to the flexibility of our proposed M-Hub, MMICT is capable of executing text generation tasks, whether based on images or videos, within a unified framework.
3. Our Method
Figure 1. Overview of MMICT. M-Hub can output both visual-guided textual features (upper left green part) and instruction-guided visual features (upper right orange part). MMICT learns from visually-guided textual features derived from demonstration examples and generates outputs based on instruction-guided visual features obtained from input queries.
@@ 3. Our Method
3.1. Multi-Modal In-Context Tuning
@@
In this study, we feed paired features from different modalities into the Multi-modal Hub (M-Hub, see Sec. 3.2) to obtain multi-modal fused features. Moreover, considering the modality gap between vision and text, we retain the visual-guided textual features as demonstrations. To elucidate the detailed formulation of MMICT, let us take video captioning as an example. As depicted in Fig. 1, given a video-instruction pair $\{v_{pred},t_{ins}\}$ (upper right orange part in Fig. 1) accompanied with other pairs $\{v_{1},t_{1},...,v_{n_{e}},t_{n_{e}}\}$ (upper left green part in Fig. 1) that are randomly selected as demonstration examples, where $\{v_{*},t_{*}\}$ denotes a video clip and the corresponding text, and $n_{e}$ is the number of demonstration examples. The model needs to predict the label $y$ (e.g., the caption text “a cat is laying down washing his face”) paired with $v_{pred}$ according to the in-context information. The frozen image encoder (e.g. EVA (Fang et al., 2022)) takes as input the video clips $\{v_{1},...,v_{n_{e}},v_{pred}\}$ and outputs the corresponding encoded visual features $\{\mathbf{Z}^{v}_{1},...,\mathbf{Z}^{v}_{n_{e}},\mathbf{Z}^{v}_{pred}\}$. We first feed these features and their paired text into M-Hub, and then pass them to a share-weight fully-connected network to extract visual-guided textual demonstration features and textual-guided visual features:
(1) $\mathbf{\hat{T}}^{c}_{k}=\text{FC}\big{(}\mathcal{G}(\mathbf{Z}_{k}^{v},t_{k})\big{)}$
(2) $\mathbf{\hat{V}}^{d}_{pred}=\text{FC}\big{(}\mathcal{G}(\mathbf{Z}_{pred}^{v},t_{ins})\big{)}$
@@
(3) $\mathbf{C}=\big{[}\mathbf{\hat{T}}^{c}_{1},\left\langle\text{EOC}\right\rangle,...,\mathbf{\hat{T}}^{c}_{n_{e}},\left\langle\text{EOC}\right\rangle,\mathbf{\hat{V}}^{d}_{pred}\big{]}$
@@
3.2. Multi-Modal Hub
Owing to the huge model size of LLMs and vision foundation models, training their parameters for multi-modal tasks proves to be challenging. As a result, many researchers endeavor to incorporate a comparatively lightweight and trainable Visual Prompt Generator (VPG) between them while maintaining their fully frozen state (Zhang et al., 2023a). Among these efforts, the BLIP-style multi-modal pre-training approach effectively connects LLMs to vision foundation models via the Q-former. However, previous works (Zhu et al., 2023; Li et al., 2023a) solely utilize this module to extract uni-modal visual features for LLMs after pre-training, thereby overlooking the benefits of multi-modal pre-training.
@@ 3.2. Multi-Modal Hub
(4) $\mathbf{P}_{h},\mathbf{R}_{h}=\text{Self-Attention}\big{(}[\mathbf{\hat{V}}^{d}_{h-1},\mathbf{\hat{T}}^{c}_{h-1}]\big{)}$
@@ 3.3. Training Objective
(5) $\mathcal{L}=-\sum_{y\in Y}\sum_{t=1}^{\left|y\right|}log\big{(}p(y_{t}|\mathbf{C},t_{ins},y_{1},...,y_{t-1})\big{)}$
@@ 3.4. Inference
(6) $\hat{y}=\text{LLM}\big{(}[\mathbf{\hat{V}}^{d}_{pred},\mathcal{T}(t_{ins})]\big{)}$
@@
4. Experiments
4.1. Implementation Details
[…]
Table 1. Instruction templates used for image captioning and video captioning.
@@ 4.1. Implementation Details
4.2. Evaluation Tasks and Datasets